OpenStudy (anonymous):

Consider the equation: x2 − y2 = 9. Find the x- and y-intercept(s) of the graph, if any exist. (If an answer does not exist, enter DNE.) x-intercept (x, y) = ____ (smaller x-value) x-intercept (x, y) =____ (larger x-value)

OpenStudy (akashdeepdeb):

For finding any intercept of any function/equation Just put x=0 for finding y intercept. Put y=0 for x-intercept....

OpenStudy (anonymous):

so would it be (-9,0) for the smaller x-value and (9,0) for the larger x value?

OpenStudy (debbieg):

If you plug in y=0, what do you get for x?

OpenStudy (anonymous):

9?

OpenStudy (debbieg):

How so? show your work.

OpenStudy (debbieg):

\( x^2 − y^2 = 9\) Put in y=0. The equation becomes...?

OpenStudy (anonymous):

x2-0=9

OpenStudy (debbieg):

Ok, simplify.... \(x^2 = 9\) Now what? How do you find x from this point?

OpenStudy (anonymous):

find the square root which is 3

OpenStudy (debbieg):

Well, yes... you take the sq. root of both sides, and what do you need to use when you do that? \(\sqrt{x^2} =???\sqrt{ 9}\) You need something here on the right hand side when you take that sq root....

OpenStudy (anonymous):

the plus and minus sign

OpenStudy (debbieg):

Exactly.... \(\sqrt{x^2} =\pm \sqrt{ 9}\) so your (two) solutions for x are?

OpenStudy (anonymous):

(-3.0) (3,0)

OpenStudy (debbieg):

yes, those are the intercepts. :)

OpenStudy (debbieg):

You see the error you were making when you said 9 and -9?

OpenStudy (anonymous):

yes i see it now, thank you so much :)

OpenStudy (debbieg):

You're welcome, happy to help. :)